OpenStudy (anonymous):

The first three terms of an arithmetic sequence are given by x, (2x -5), and 8.6.

OpenStudy (anonymous):

What is the first term and the common difference?

OpenStudy (inkyvoyd):

Well the differnce is "common" for a reason. Set the difference equal, solve for x, and you have your answers.

OpenStudy (saifoo.khan):

a, b, c common diff = c-b or b-a

OpenStudy (anonymous):

What is the reasoning for (2x - 5) - x = 8.6 - (2x - 5)

OpenStudy (anonymous):

Why do we make them equal

OpenStudy (inkyvoyd):

FOO FOO STEAL MAH MEDAL

OpenStudy (inkyvoyd):

Because we know in an arithmeic series, that each term has the same difference as the preveious

OpenStudy (inkyvoyd):

It's actually by definition.

OpenStudy (anonymous):

I understand that they equal, d is equal to d but why can't I just go 2x - 5 - x = the answer?

OpenStudy (inkyvoyd):

Because. That doesn't simplify to a constant.

OpenStudy (anonymous):

Is it because you require an equation to solve the problem and not an expression?

OpenStudy (inkyvoyd):

Yes.

OpenStudy (saifoo.khan):

it's correct.

OpenStudy (inkyvoyd):

It's because you get two expressions, actually. Those two expressions are equal, but not equivalent.

OpenStudy (anonymous):

But 2x - 5 - x would be x - 5 which is just an expression right? but we need an equation to solve the problem? I am trying to understand the philosophy behind how this works.

OpenStudy (inkyvoyd):

(in the sense that x+3+8 is equivalent to x+2+9)

OpenStudy (saifoo.khan):

if you think, a-b = b-c both will give you the same value.

OpenStudy (inkyvoyd):

Well. We actually get two expressions.

OpenStudy (inkyvoyd):

a-b = b-c, as foo foo said.

OpenStudy (anonymous):

We do but the expressions equal eachother right?

OpenStudy (saifoo.khan):

yes.

OpenStudy (saifoo.khan):

we will get a value for 'x' then.

OpenStudy (inkyvoyd):

a-b is an expression, and b-c is also an expression. Each of these is the comon difference.

OpenStudy (anonymous):

(2x - 5) - x = x - 5 you CAN'T make it equal to zero right?

OpenStudy (inkyvoyd):

Wat do you mean, equal zero?

OpenStudy (saifoo.khan):

No, not equal to zero

OpenStudy (anonymous):

Like if you went 2x -5 - x = x -5 then took x - 5 = 0, x = 5

OpenStudy (inkyvoyd):

Because x-5 is the common difference.

OpenStudy (inkyvoyd):

not zero; the common diference

OpenStudy (saifoo.khan):

(2x - 5) - x = 8.6 - (2x - 5) x = 6.2

OpenStudy (anonymous):

I see But making a quadratic equal 0 works because it when you make a quadratic equal to 0 it finds the x intercepts right?

OpenStudy (inkyvoyd):

We know a value x-5 is the common difference, and also a value 2x-5-x, which is also the common difference. We can set them equal because we know the represent the same value (the common difference)

OpenStudy (inkyvoyd):

Well, you set it equal to zero

OpenStudy (inkyvoyd):

because in a quadratic function y=ax^2+bx+c, you are looking for the x-value when y is equal to zero (in other words, the x intercept)

OpenStudy (inkyvoyd):

the point (x,0) is the x intercept.

OpenStudy (anonymous):

Another question, adding a negative is the same as minusing a positive and minusing a postive is the same as adding a negative right?

OpenStudy (inkyvoyd):

There are two unknowns in a quadratic function.

OpenStudy (anonymous):

Thanks

OpenStudy (inkyvoyd):

When we search for the x-intercept, we know that y=0, because the x intercept is on the x-axis and has no vertical height.

OpenStudy (inkyvoyd):

Another question, adding a negative is the same as minusing a positive and minusing a postive is the same as adding a negative right? - those are properties the directly follow from our definition of negative numbers and subtraction, as well as addition.